OpenStudy (anonymous):

what percent of 40 is 20?

sam (.sam.):

\[\frac{x}{100}*40=20\] x=50

OpenStudy (anonymous):

20% of what number is 40 ?

OpenStudy (anonymous):

50 percent

OpenStudy (anonymous):

i knew that thx

OpenStudy (anonymous):

8% of 125 is what number?

OpenStudy (anonymous):

percent always means divide by 100 of always means multiply is always means equals and what number always means X

OpenStudy (anonymous):

WHAT

OpenStudy (anonymous):

8 % of 125 is what number? 8 / 100 * 125 = X

OpenStudy (anonymous):

ooh 15% of what number is 24?

OpenStudy (anonymous):

what percent of 40 is 20? X / 100 * 40 = 20

OpenStudy (anonymous):

it's not odd, it's how you translate English into algebraic notations

OpenStudy (anonymous):

so my answer is 12.5

OpenStudy (anonymous):

I dont know about answers... thats for calculators... what equation did you use?

OpenStudy (anonymous):

divide

OpenStudy (anonymous):

normal calculator

OpenStudy (anonymous):

well, ok then I trust them, just so long as you press the right buttons

OpenStudy (anonymous):

ooh 15% of what number is 24 ?

OpenStudy (anonymous):

translate it

OpenStudy (anonymous):

percent always means divide by 100 of always means multiply is always means equals and what number always means X

OpenStudy (anonymous):

\[100\div15=x\]

OpenStudy (anonymous):

15/100 is 15%

OpenStudy (anonymous):

15% of what number is 24? % means /100 15/100 of what number is 24? of means multiply 15/100 * what number is 24? what number means X 15/100 * X is 24? is means = 15/100 * X = 24

OpenStudy (anonymous):

i give up i might as well just get a "F" on my report card

OpenStudy (anonymous):

close this one and repost it.... maybe someone else can explain it better than I sorry if I confused you